On January 1, 2021, the Marjlee Company began construction of an office building to be used as its corporate headquarters. The building was completed early in 2022. Construction expenditures for 2021, which were incurred evenly throughout the year, totaled $8,700,000. Marjlee had the following debt obligations which were outstanding during all of 2021: Construction loan, 9% $ 2,175,000 Long-term note, 8% 2,900,000 Long-term note, 5% 5,800,000 Required: Calculate the amount of interest capitalized in 2021 for the building using the specific interest method.
Answer #1

Interest capitalized in 2021 for the building using the specific interest method = $326,250

Explanation

Average accumulated expenditures = $8,700,000 / 2 = $4,350,000
Weighted average rate of all other debt
Amount ($) x Rate of interest = Weighted average amount ($)
        2,900,000.00 x 8% =                             232,000.00
        5,800,000.00 x 5% =                             290,000.00
     8,700,000.00                           522,000.00
Weighted average rate = 522,000 / 8,700,000 = 6%
Interest capitalized:
Amount ($) x Rate of interest = Interest capitalized ($)
        2,175,000.00 x 9% =                             195,750.00
        2,175,000.00 x 6% =                             130,500.00
TOTAL 326,250.00